OpenStudy (anonymous):

The police car with its 300-Hz siren is moving away from the warehouse at a speed of 16.0 m/s . What frequency does the driver of the police car hear reflected from the warehouse? Use 344m/s for the speed of sound in air.

OpenStudy (anonymous):

This is Doppler effect problem. Remember this one?

OpenStudy (anonymous):

The equation? fL =v+vL/lambda

OpenStudy (anonymous):

It's a little bit different from the one i know. \[f_r=\frac{v\pm v_r}{v\pm v_s}f_s\] I'm sorry but I have to leave now.

OpenStudy (anonymous):

Its ok. And that equation looks familiar so yours is probably right

OpenStudy (anonymous):

Someone Please Help!

OpenStudy (anonymous):

286 Hz

OpenStudy (anonymous):

Is it exactly 286 hertz? Because I only have one more attempt at this problem

OpenStudy (anonymous):

Actually, I tried that earlier and it was wrong

OpenStudy (anonymous):

Maybe with the reflection off the warehouse, it is less, because the warehouse "percieves" a lower frequency and reflects it back to the car so maybe this is the answer: (344-2*(16))*300/344 = 272 Does this seem reasonable?

OpenStudy (anonymous):

Is there any other possible answers?

OpenStudy (anonymous):

I only have one try left so I want to make sure

OpenStudy (anonymous):

That is the best I can think of without having to review a textbook on Doppler effect. It seems the frequency is reduced twice, once traveling to the warehouse and again with the reflected soundwave. I think it is right.

OpenStudy (anonymous):

ok thanks. I'll try it

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

Thanks!!! it worked. Do you think you can help me with another problem?

OpenStudy (anonymous):

Okay, sure. What is it?

OpenStudy (anonymous):

I'll close this and open a new question